The rellaw pedaling the sieycle with r660omm wberh at 30u/h braies and sp in 45m. What is Hs deceration? How many cond, withi? How mary e. our ansswers fromd& o by substauting the results in a ghird equation
Answer #1

initial speed = 30km/hr = 8.33 m/sec, final speed = 0, distance travelled = 4.5 m to find a = ?

v2 = u2 + 2*a*S => 0 = 8.332 + 2*a*4.5 => a = -7.72 m/s2

total time taken v = u + a*t   => 0 = 8.33 - 7.72t = 1.08 sec